TCCON > Sites > Caltech/Pasadena > Instrument History > Alignment History

Alignment History

December 15, 2011

Interferometer Alignment

Before alignment

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A 8.4V -8.84V -320mV -8.32V 0.93
Front B 8.0V -9.24V -1.28V -8.56V 0.926
Back A 7.92V -8.88V -320mV -7.96V 0.835
Back B 7.6V -9.24V -1.4V -8.56V 0.892

First alignment of interferometer

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A 8.48V -8.8V -120mV -8.28V 0.92
Front B 8.24V -9.24V -880mV -8.8V 0.94
Back A 8.24V -8.84V -240mV -8.04V 0.86
Back B 8.32V -9.2V -1.20V -8.44V 0.91

Ghost Reduction

GPR table

Back to top

April 30, 2012

Interferometer Alignment

Check of interferometer alignment - no adjustment seems necessary. These are still better than the latest Lamont values.

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A 10.75V -9.25V 1.6875V -9.0625V 0.966
Front B 9.875V -9.4375V 750mV -9.125V 0.940
Back A 10.00V -9.1875V 1.4375V -8.5625V 0.889
Back B 9.375V -9.4375V 562.5mV -8.8125V 0.882

IR Alignment

Using a 1.3 mm aperture results in a fairly poor ILS. Note, used 8-fold zero-filling in OPUS. Using 1-fold zero-filling caused an undersampling error in LINEFIT:

me_phs_20120430_APT1.3mm.png

It looks better when using a 1.0 mm aperture, but still not great:

me_phs_20120430_APT1mm.png

Back to top

May 1, 2012

After IR Alignment

This is using the "spare" detector with the SMC connector on the back: not the InGaAs detector that came with this interferometer.

Much better! The following is with 1.3 mm APT:

me_phase_20120501_apt1.3.png

And with 1.0 mm APT:

me_phase_20120501_apt1.png

May 2, 2012

Ghost Reduction 

  Spectrum
Name
Parent
Wavenumber
Parent
Amplitude
Ghost
Wavenumber
Ghost
Amplitude
Scanning
Frequency
Resolution GPR Notes
Before adjustment ci20120502-07-ECL00gpr.00 5912.389 0.26671 11804.586 0.0004147 7.5 kHz 2 cm-1 0.0015548

Not the laser sampling ghost! This is the double-frequency ghost!

Before adjustment ci20120502-07-ECL00gpr.00 5912.389 0.26671 9830.300 0.00000705 7.5 kHz 2 cm-1 2.64e-5 Laser-sampling ghost. Small ghost visible.
After adjustment ci20120503-07-ECL00.gpr.03 5912.234 0.26720 9830.300 0.00000670 7.5 kHz 2 cm-1 2.51e-5 Laser-sampling ghost. No ghost visible. Measuring top of noise, so overestimating the GPR.

ghost_reduction.bmp

September 5, 2012

Interferometer Alignment

Laser died, installing new laser.

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A 5.325V -8.925V -3.5V  -8.825V  0.96
Front B 5.1V -9.275V -4.1V  -9.2V  0.97
Back A 5.125V -8.95V  -3.575V  -8.7V  0.91
Back B 5.025V -9.275V  -4.125V  -9.15V  0.95

After fastening on the cover

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A  6.5V -8.925V -2.275V -8.775V 0.96
Front B  5.3V -9.3V -3.875V -9.175V 0.95
Back A  6.2V -8.925V -2.375V -8.6V 0.91
Back B  5.0V -9.275V -4.05V -9.025V 0.91

Ghost Minimization

Ghost:Parent Ratios

  Spectrum
Name
Parent
Wavenumber
Parent
Amplitude
Ghost
Wavenumber
Ghost
Amplitude
Scanning
Frequency
Resolution GPR Notes
Before adjustment if20120905-07-ECL00gpr.00 5914.7 0.255 9893.804 0.0000546 7.5 kHz 1 cm-1   Laser-sampling ghost. Small ghost visible.
Before adjustment ci20120905-07-ECL00gpr.00 5912.441 0.25785 9893.086 0.00003171 7.5 kHz 2 cm-1  0.00012297847 Laser-sampling ghost. Ghost visible
After adjustment ci20120905-07-ECL00gpr.02 5912.724 0.257170 9891.042 0.00000137 7.5 kHz 4 cm-1 0.00000533  

 

September 7, 2012

Laser has been turned 180 degrees. Start again.

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A  7.925  -8.875V  -650mV  -8.575V  .93
Front B  7.3  -9.275V -1.3V  -8.6V  .84
Back A  7.75  -8.9V  -650mV  -8.4V  .89
Back B  7.15  -9.275V  -1.35V  -8.5  .82

This looks terrible. We can't seem to get the ILS symmetric:

 me_phase_20120907.png

September 12, 2012

Put in Jean-Francois's Spectra Physics laser. Start again.

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A   -9.0 4.3125 -8.75 0.96
Front B   -9.375 2.75 -8.5 0.87
Back A   -9.125 4.125 -8.625 0.93
Back B   -9.4375 2.5625 -8.375 0.88

 

Ghost Minimization
Edit section

Ghost:Parent Ratios

  Spectrum
Name
Parent
Wavenumber
Parent
Amplitude
Ghost
Wavenumber
Ghost
Amplitude
Scanning
Frequency
Resolution GPR Notes
Before adjustment ci20120912-07-ECL00gpr.01  5912.859 0.205669 9885.401 0.00004347 7.5 kHz 1 cm-1  2.1e-4 Laser-sampling ghost. Large ghost visible.
After adjustment ci20120912-07-ECL00gpr.0  5912.487 0.20581 9885.101 0.00000147 7.5 kHz 4 cm-1  7.1e-6  

 

September 14, 2012

  • Unable to align the instrument to our satisfaction. There is clear, significant asymmetry in the line shape.
  • We've ruled out:
    • OPUS-IPP vs OPUS (when processing the interferograms, the spectra look the same through both codes)
    • Reseating the beamsplitter (no change)
    • Changing the laser (no real change)
    • Changing the InGaAs detector (no change)
  • Things to try:
    • Replace laser sampling board
    • Replace beamsplitter
    • Collect data with IFSloop instead of OPUS

November 2, 2012

  • Adjusted laser detector assembly positions: installed lock washers underneath to give more DOF.
Location Laser Zero Max Min Modulation Efficiency
Front A -8.95 5.725 -8.425 0.93
Front B -9.25 2.75 -8.45 0.88
Back A -8.875 5.625 -8.35 0.93
Back B -9.225 2.25 -8.325 0.85

ILS_20120828_20121102.png

Back to top

July 10, 2015

Results from prior to today. Note that the scanner/encoder was swapped April 20, 2015.

ILS_201506.jpg

Perhaps no need for IR alignment? Though at 75 cm OPD, there may be overmodulation, which would indicate a shear error:

Laser Alignment

  • Alignment after the scanner/motor assembly was replaced, which was causing noise bursts to create ghosts in the system. This is likely (hopefully) the reason why the alignment was so unstable before.
  • Adjusted the fixed cube corner and MOPD.
Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A 10.6V -9.0V 2.475V -7.825V  0.814
Front B 10.37V -9.3V -875mV -8.65V  0.857
Back A   -9.0V 2.575V -7.7V  0.798
Back B   -9.3V -950mV -8.475V  0.82
After ZPD Adjustment            
Front A    -9.0V  3.175V  -8.025  0.85
Front B    -9.3V  -0.425V  -8.85V  0.90
Back A    -9.0V  2.75V  -7.825  0.81
Back B    -9.3V  -0.725V  -8.6V  0.85
After MOPD Adjustment            
Back  A     3.125V  -8.025V  0.85
Back  B     -0.550V  -8.8V  0.89
Front  A      3.225V  -8.025V  0.85
Front  B      -0.425V  -8.8V  0.89

 

IR Alignment

before_pink_after_blue_alignment_caltech_ifs9_20150710.bmp

Solar Alignment

  • No sun available when we finished.

Ghosts Minimization

  Spectrum
Name
Parent
Wavenumber
Parent
Amplitude
Ghost
Wavenumber
Ghost
Amplitude
Scanning
Frequency
Resolution GPR Notes
Before adjustment  ci20150710laaaag.002 5912  0.052 9822.8  0.0000024 10 kHz 4 cm-1 4.6e-5  Visible
After adjustment  ci20150710laaaag.004 5912.724  0.051 9891.042  0.000001 10 kHz 4 cm-1 1.9e-5  Can't see it.

Results

Looks like we made the shear error worse after this alignment:

ILS_20150717_45cmOPDTCCON.jpg

Back to top

July 24, 2015

We centered the fringes of our "alignment kit" (external HeNe laser on a camera, lens, prism, speckle reducer, and camera) using the two flat mirrors directing the beam from the first field stop into the interferometer (the IR Adjustment Screws) and compensating with IR Adjustment Screw Compensating Mirror.  Then, we measured the first set of Mod. Eff. values.  Next, we adjusted the laser entrance mirror (the three precision screws on the mount next to the beamsplitter enclosure - closest to the LASB detector - the MOPD Laser Adjustment Screws).  After several iterations, we recorded the "After Adjustment" values below and very last spectrum of the day.

Location Laser Peak-to-Peak Zero Max Min Modulation Efficiency
Front A 10.5 -8.95V 2.225V -8.275V  0.89
Front B 7.9 -9.3V -1.05V -8.95V  0.92
Back A 10.4 -9.0V 2.075V -8.325V  0.89
Back B 7.418 -9.25V -1.357V -8.775V  0.89
After Adjustment            
Front A  10.7  -8.975V 2.425V  -8.275  0.88
Front B  8.175  -9.275V  -825mV  -9.0V  0.94
Back A  10.125  -8.975V  2.125V  -8.0V  0.84
Back B  7.725  -9.275V  -1.1V  -8.825V  0.90

 

before_pink_after_blue_worst_green_alignment_caltech_ifs9_20150724.bmp 

The pink trace is the original line shape before alignment on July 24, the green is after the IR alignment (but not the met. laser alignment), and the blue is after the laser alignment (under vacuum). Note that you should always measure at 10 kHz, and not 7.5 kHz, as the 7.5 kHz speed is noisy.

The traces above are scaled using the OPUS calculator option, they are not rescaled using the autoscaler, so the zero is the same for all traces. From Jean-Francois:

You'll remember that I always curse OPUS because the plot auto-scale does not preserve the zero.  Today, Keeyoon showed me a nice solution to this problem.  One can click on the calculator icon and use that to multiply a spectrum by any factor.  This allows us to line up the continuum while keeping the zero fixed.  Then one can see if the lines get deeper during instrument alignment.

Indeed, the alignment is now much better.

ILS_20150728_45cmOPDTCCON.jpg

Back to top

Tag page

Files 2

FileSizeDateAttached by 
 ci20150710laaaaa.012
Final lamp spectrum after alignment on July 10, 2015
52.95 MB13:38, 10 Jul 2015dwunchActions
 ci20150727laaaaa.001
Final lamp spectrum after alignment on July 24, 2015
26.48 MB07:18, 27 Jul 2015dwunchActions